A company is financed for 30% of its capital with debt and the rest with equity. debt is risk free, so it’s beta is B=0 and its cost is risk free rate rf. Equity has a beta of Be=2

1. What is the firms asset beta.

2. Asume CAPM is correct. What is the cost of capital for the firm? Assume risk free interest of 5% and market risk premium of 6%
Answer #1

Weight of Debt = wD = 0.3, cost of Debt = rf, beta of Debt = βD = 0

Weight of Equity = wD = 0.7, Beta of Equity = βE = 2

1. Asset beta of the firm is calculated using the formula:

βA = wDD + wEE = 0.3*0 + 0.7*2 = 0 + 1.4 = 1.4

Answer -> Firm's asset beta = βA = 1.4

2. Risk-free rate = rF = 5%

Market return = rM

Market risk premium = (rM - rF) = 6%

Cost of capital of the firm = rA

Cost of capital of the firm can be calculated using CAPM

rA = rF + βA*(rM - rF)

rA = 5% + 1.4*6% = 13.4%

Answer -> cost of capital for the firm = 13.4%